33S NUCLEAR QUADRUPOLE COUPLING IN THIIRANE, METHYLTHIIRANE, 2,2-DIMETHYLTHIIRANE, AND ANTI-2,3-DIMETHYLTHIIRANE: HYPERFINE STRUCTURE IN THE ROTATIONAL SPECTRA AND AB-INITIO CALCULATIONS

\underline{JENS-UWE GRABOW}, HELMUT DREIZLER, Institut für Physikalische Chemie, Christian-Albrechts-Universität, Olshausenstraße 40, D-24098 Kiel, Germany; BARBARA KIRCHNER, HANSPETER HUBER, GEROLD STEINEBRUNNER, Institut für Physikalische Chemie, Universität Basel, Klingelbergstraße 80, CH-4056 Basel, Switzerland; FELIPE LORENZO, JOSE L. ALONSO, Departamento de Quimica, Facultad de Ciencias, Universidad de Valladolid, Prado de la Magdalena, s/n, SP-47005 Valladolid, Spain; ILONA MERKE, Institut für Physikalische Chemie, Rheinisch-Westfälische Technische Hochschule, Templergraben 59, D-52062 Aachen, Germany; MASAO ONDA, Department of Chemistry, Faculty of Science and Technology, Sophia University, Kioicho 7-1, Chiyoda-ku, Tokyo 102, Japan.

We present an investigation of the nuclear quadrupole coupling in a series of substituted thiiranes along with thiirane itself: The rotational spectra and the quadrupole coupling hyperfine structures of ^{33}S thiirane, ^{33}S methylthiirane, ^{33}S 2,2dimethylthiirane, and ^{33}S anti-2,3-dimethylthiirane in the range of 7 to 22 GHz are reported. Quantum chemical calculations on the MP4(QSD) level with basis sets of high local quality were performed to determine the nuclear quadrupole coupling tensor of ^{33}S. The analysis of the nuclear quadrupole coupling in the rotational spectra provided experimental information on the tensors.